# Agent-as-Judge for Factual Summarization of Long Narratives

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated near-human performance in summarization tasks based on traditional metrics such as ROUGE and BERTScore. However, these metrics do not adequately capture critical aspects of summarization quality, such as factual accuracy, particularly for long narratives (>100K tokens). Recent advances, such as LLM-as-a-Judge, address the limitations of metrics based on lexical similarity but still exhibit factual inconsistencies, especially in understanding character relationships and states. In this work, we introduce NarrativeFactScore (NFS), the first “Agent-as-a-Judge” framework that evaluates and refines factuality in narrative summarization. By leveraging a Character Knowledge Graph (CKG) extracted from input narrative, NarrativeFactScore evaluates the factuality and provides actionable guidance for refinement, such as identifying missing or erroneous facts. Our experimental results demonstrate that constructing the CKG enables reasoning with 1/3 of the factuality computation used in prior approach, and achieve three times higher correlation with human judgments. Furthermore, refinement with actionable guidance improves the quality of the summary. [^1] https://github.com/YeonseokJeong/NarrativeFactScore
@@
Yeonseok Jeong1, Minsoo Kim1, Seung-won Hwang2thanks: Corresponding Author, Byung-Hak Kim3 IPAI, Seoul National University1, Seoul National University2, CJ Group3 {jys3136, minsoo9574, seungwonh}@snu.ac.kr bhak.kim@cj.net
@@ 1 Introduction
However, directly using LLM to evaluate factuality has two limitations. First, as demonstrated by Kim et al. (2024), the LLM judge often fails to accurately assess factuality in narratives that require indirect reasoning, such as understanding character relationships or states. For example, in Figure 1, although Sauron is pursuing Frodo in order to obtain the Ring in The Lord of the Rings, the LLM judge inaccurately evaluates the factuality of a summary which incorrectly reports that “Sauron is pursuing Gandalf”. This limitation stems from the inability of the LLM judge to consistently track and reason about character relationships, highlighting the need to maintain structured and consistent CKG.
Second, the LLM judge evaluates factuality as a score with no interpretability, whereas ROUGE Lin (2004) and BERTScore Zhang et al. (2019) enable to interpret why a given score was assigned and further facilitates the generation of actionable feedback. Desirably, evaluation metrics for summary can provide feedback, when the score is low, to explain why it is incorrect and suggest how to improve.
@@
To improve the interpretability of the metric, NarrativeFactScore also provides feedback for interpretation and refinement when the summary is incorrect. For each statement in the summary, our metric retrieves relevant scenes and character relationships from our CKG to calculate a factuality score. Based on the retrieved evidence, ours evaluates each statement and generates feedback identifying discrepancies between claims and supporting evidence. Since our metric operates autonomously, it is more cost-effective and faster than Human-as-a-Judge. In addition, it offers feedback for low scores, which makes it more reliable than LLM-as-a-Judge metrics. Recognizing the causes of low scores also contributes to generating more accurate summaries through agent-based refinement.
Using NarrativeFactScore provides two key advantages for long narrative summarization. First, it offers a labor-efficient and fast metric that also approximates human evaluation when evaluating the factuality of summaries. Our metric demonstrates a statistically strong correlation with human evaluation, and a test for differences between human evaluation and our metric yielded statistically significant results, with the p-value falling below 0.05. Second, since it provides feedback on factually incorrect parts, agent-based refinement can improve summarization performance. We show that agent-based refinement improves factuality (+14.03), ROUGE (+2.05), and BERTScore (+0.13) on MovieSum Saxena and Keller (2024a), a movie script summarization dataset, and also improves factuality (+12.26), ROUGE (+2.47), and BERTScore (+0.21) on MENSA Saxena and Keller (2024b), a movie scene saliency dataset.
@@ 2.1 Long Narrative Summarization
Summarizing long narratives Saxena and Keller (2024a, b) is challenging due to the high computational and memory demands required by transformer-based models. In prior work Pilault et al. (2020); Li et al. (2021); Wu et al. (2021); Chang et al. (2024), a method called hierarchical merging was introduced, where individual chunks of the narrative are summarized separately and then combined to form a coherent final summary. Although this method preserves the logical structure of the narrative, hallucinations remain a frequent challenge, especially when capturing global information such as character relationships. Thus, our focus is on improving the factuality of the summaries.
@@ 2.2 Character Knowledge Graph (CKG)
Since characters are integral to narrative Gurung and Lapata (2024), prior work has aimed to construct a graph to easily utilize them. In narrative texts, CKG shows the unidirectional relationship between a subject and an object character. This process is similar to creating a triple (subject-predicate-object) list in knowledge graph construction Chen et al. (2020). Andrus et al. (2022) utilized the OpenIE system Angeli et al. (2015) for story completion and question-answering tasks, integrating it with GPT-3 Brown et al. (2020) to enhance its effectiveness. Alternatively, a recent method Zhao et al. (2024) that assembles CKG directly using LLMs is more robust approach, as it better captures the nuanced and complex relationships. Our distinction lies not only in constructing CKGs but also in utilizing them to measure and enhance factuality.
@@ 3.1 CKG Extraction
Given a narrative represented as a collection of scenes $\mathcal{N}=\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{m}\}$, where $m$ denotes the number of scenes, the goal is to extract a graph $G$ that encapsulates character relationships. Each scene $\mathcal{S}_{i}\ (1\leq i\leq m)$ is processed individually to extract relation triples (subject-predicate-object) using GPT-4o-mini OpenAI (2023), as detailed in Section D.1. The extracted triples are used to initialize the nodes and determine the edges based on the main relationships between the nodes, forming the final CKG $G$ through the following two steps.
@@
First, to maintain consistency in character identification, we construct a names graph $G_{name}$, consolidating aliases or variations in names in scenes. Our framework processes each scene in turn, extracts all character names, and determines several names refer to same character based on the context using LLM. For example, in The Lord of the Rings, ‘Frodo’ and ‘Frodo Baggins’ are recognized as the same character. As illustrated by ‘Frodo / Frodo Baggins’ in Figure 3(a), each name variation is a node. [^2] In practice, an undirected edges is added between nodes that refer to the same character. This step ensures an accurate capture of relationships, even when names vary across scenes. The CKG is initialized using names from the names graph.
Second, to preserve the consistency of relationships, we sample extracted triples multiple times Wang et al. (2023); Brown et al. (2024) and select frequent ones as the final edges. Let the node set $V$ be the set of all character in the names graph:
@@
Only triples with named entities as subjects and objects are used; if an object is missing, a self-loop is added to represent the state of character. We then define the edge set $E$ of our CKG as
@@
For triples with the same subject and object, frequent predicates capture temporal changes as directed edges. For example, since the early scenes show that ‘Frodo’ fears ‘Sauron’ and the later scenes show that he resists ‘Sauron’, the CKG in Figure 3(a) displays two distinct relationships. The process of deciding edges is repeated to construct a CKG that can effectively evaluate the factuality of summaries.
@@ 3.2 NarrativeFactScore Calculation
We invent a new metric to guide agentic evaluation, unlike existing factuality metrics Min et al. (2023); Bishop et al. (2024) that do not provide evidence or feedback for their scores, by considering events in the input story superficially but overlooking relational information about characters. Our metric addresses these limitations by incorporating character relationship graphs and providing detailed feedback. To calculate the factuality of the narrative summary, we first generate an initial summary $Z$ using the prompt described in Section D.2.
To evaluate the factuality of the initial summary $Z$, we decompose it into smaller verifiable units, similar to the approach used in Min et al. (2023). Using the prompt in Section D.3, each sentence in the initial summary $Z$ is divided into a list of atomic facts $A=\{a_{1},a_{2},\ldots,a_{z}\}$.
To evaluate each atomic fact $a_{k}$, we need the scene and information about the characters that appear in the atomic fact. First, we retrieve the most relevant scene $\mathcal{S}_{i}$ within the narrative $\mathcal{N}$, by using the BGE-M3 Chen et al. (2024). Second, we also retrieve the subgraph $g$ from the linearized CKG $G$, as illustrated in Figure 3(b), which contains triples involving the characters mentioned in $a_{k}$.
@@
where $\mathbbm{1}$ is the indicator function, yielding 1 if the atomic fact $a_{i}$ is factual and 0 otherwise. This evaluation is carried out using the prompt detailed in Section D.4, which produces 1 if the atomic fact is accurate. If the atomic fact is determined to be inaccurate, then feedback $f_{k}$ is provided on how to correct it. Finally, the NarrativeFactScore is calculated as the proportion of atomic facts that are found to be factual, defined by the following equation:
@@ 3.3 Agent-based Fact Refinement
The new metric leveraging consistent CKG enables LLM agent to guide refinement by using feedback from the evaluation. This process involves three key inputs: original narrative to provide global context, the initial summary that requires modification, and the feedback detailing the inaccuracies and reasons for those errors. Using a prompt that incorporates these inputs in Section D.5, LLM refines the initial summary by correcting only the factually inaccurate parts provided from feedback, then generates the improved summary $y$.
@@ 4.1 Implementation Details
To generate the initial summary $Z$, we adopt hierarchical merging Chang et al. (2024) that ensures the logical structure of the narrative is preserved. The narrative is first divided into chunks $C_{i}$ where each chunk is formed incrementally by adding scenes until a predefined context size444we set predefined context size of a chunk to 1024. is reached. Once this limit is exceeded, a new chunk begins, resulting in a sequence of chunks $\mathcal{C}=\{C_{1},C_{2},...,C_{n}\}$. Each chunk $C_{i}$ is then independently summarized using the prompt specified in Section D.2, and the resulting chunk summaries are sequentially merged to produce the initial summary $Z$.
@@ 4.3 Correlation with Human Factuality Scores
To evaluate whether the NarrativeFactScore we proposed correlates effectively with human factuality, we conducted a series of experiments. For this purpose, we used STORYSUMM  Subbiah et al. (2024) and FABLES Kim et al. (2024), which include multiple summaries generated by LLM for each narrative. These summaries were then evaluated by human annotators based on their factual accuracy.
@@ 4.4 Summarization Performance Evaluation
We evaluated our framework on the MENSA Saxena and Keller (2024b) and MovieSum Saxena and Keller (2024a) datasets. MENSA aligns movie scenes with Wikipedia summaries and MovieSum pairs screenplays with summaries. We use the full test sets: 50 samples from MENSA and 200 from MovieSum.
@@ 5 Analysis
5.1 Analysis for Baseline Metrics
Strength Spearman ($\rho$) KENDALL ($\tau$)
Very weak correlation 0.00~0.15 0.00~0.10
Weak correlation 0.15~0.30 0.10~0.20
Moderate correlation 0.30~0.43 0.20~0.30
Strong correlation 0.43~1.00 0.30~1.00
Table 4: Correlation strength based on coefficients.
[…]
@@ 5.2 How Consistently Does Ours Capture Character Relationships?
To effectively evaluate factuality and improve summary, it is necessary to generate an accurate and consistent CKG. According to Kim et al. (2024); Zhao et al. (2024), the “naive extract” approach, where an LLM extracts character relationships in one step, often fails to consistently capture some relationships. Thus, our objective is to verify whether our approach can generate a consistent CKG. Conan Zhao et al. (2024) provides ground truth annotation of character relationships within narratives. To evaluate whether the generated relation is semantically similar to this ground truth, we measure the BERTScore Devlin et al. (2019).
As shown in Table 5, our method generates CKG that are closely similar to ground truth. Although the “naive extract” achieves 86.26, it occasionally produces incorrect relationships. In contrast, by reasoning about relationships scene by scene and aggregating them, our method chooses more accurate relationships and constructs a consistent CKG.
@@ 5.3 Challenging Set
We aim to evaluate whether our metric can provide feedback necessary to improve factuality in recent narratives. Although LLM-based metrics provide accurate factuality feedback for narratives within their pretraining data, they often fail for narratives outside of their training corpus. However, our metric provides accurate feedback by evaluating summaries based on narrative story and character relationships rather than relying on parametric knowledge alone. Therefore, we define a challenging set of works published after the knowledge cutoff date of our LLM to verify whether our metric improves factuality through its feedback.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated near-human performance in summarization tasks based on traditional metrics such as ROUGE and BERTScore. However, these metrics do not adequately capture critical aspects of summarization quality, such as factual accuracy, particularly for long narratives (>100K tokens). Recent advances, such as LLM-as-a-Judge, address the limitations of metrics based on lexical similarity but still exhibit factual inconsistencies, especially in understanding character relationships and states. In this work, we introduce NarrativeFactScore (NFS), the first “Agent-as-a-Judge” framework that evaluates and refines factuality in narrative summarization. By leveraging a Character Knowledge Graph (CKG) extracted from input narrative, NarrativeFactScore evaluates the factuality and provides actionable guidance for refinement, such as identifying missing or erroneous facts. Our experimental results demonstrate that constructing the CKG enables reasoning with 1/3 of the factuality computation used in the prior approach, and achieve three times higher correlation with human judgments. Furthermore, refinement with actionable guidance improves the quality of the summary. [^1] https://github.com/YeonseokJeong/NarrativeFactScore
@@
Yeonseok Jeong1, Minsoo Kim1, Seung-won Hwang2thanks: Corresponding Author, Byung-Hak Kim3 IPAI, Seoul National University1, Seoul National University2, Hyundai Card3 {jys3136, minsoo9574, seungwonh}@snu.ac.kr byunghak.kim@hcs.com
@@ 1 Introduction
However, directly using LLM to evaluate factuality has two limitations. First, as demonstrated by Kim et al. (2024), the LLM judge often fails to accurately assess factuality in narratives that require indirect reasoning, such as understanding character relationships or states. For example, in Figure 1, although Sauron is pursuing Frodo in order to obtain the Ring in The Lord of the Rings, the LLM judge inaccurately evaluates the factuality of a summary which incorrectly reports that “Sauron is pursuing Gandalf”. This limitation stems from the inability of the LLM judge to consistently track and reason about character relationships. To address this, we introduce the CKG, a structured representation of characters and their relationships that must be maintained consistently.
Second, the LLM judge outputs only a single score with limited interpretability, which makes it less reliable and difficult to identify what needs to be improved. Desirably, evaluation metrics for summarization can provide feedback, when the score is low, to explain why it is incorrect and suggest how to improve.
@@
To improve the interpretability of the metric, NarrativeFactScore also provides feedback for interpretation and refinement when the summary is incorrect. For each statement in the summary, our metric retrieves relevant scenes and character relationships from our CKG to calculate a factuality score. Based on the retrieved evidence, ours evaluates each statement and generates feedback identifying discrepancies between claims and supporting evidence. Since our metric operates autonomously, it is more cost-effective and faster than Human-as-a-Judge. In addition, it offers feedback for low scores, which makes it more interpretable than LLM-as-a-Judge metrics. Recognizing the causes of low scores also contributes to generating more accurate summaries through agent-based refinement.
Using NarrativeFactScore provides two key advantages for long narrative summarization. First, it offers a labor-efficient and fast metric that also approximates human evaluation when evaluating the factuality of summaries. Our metric demonstrates a statistically strong correlation with human evaluation, and a test for differences between human evaluation and our metric yielded statistically significant results, with the p-value falling below 0.05. Second, by providing feedback on factually incorrect parts, it can improve summarization performance. We show that agent-based refinement improves factuality (+14.03), ROUGE (+2.05), and BERTScore (+0.13) on MovieSum Saxena and Keller (2024a), a movie script summarization dataset, and also improves factuality (+12.26), ROUGE (+2.47), and BERTScore (+0.21) on MENSA Saxena and Keller (2024b), a movie scene saliency dataset.
@@ 2.1 Long Narrative Summarization
Summarizing long narratives Saxena and Keller (2024a, b) is challenging due to the high computational and memory demands required by transformer-based models. In prior work Pilault et al. (2020); Li et al. (2021); Wu et al. (2021); Chang et al. (2024), a method called hierarchical merging was introduced, where individual chunks of the narrative are summarized separately and then combined to form a coherent final summary. Other segmentation-aware strategies have also been explored Moro and Ragazzi (2022, 2023); Zhang et al. (2022). But such chunk-based methods can group unrelated content into the same segment. Although these methods preserve the logical structure of the narrative, hallucinations remain a frequent challenge, especially when capturing global information such as character relationships. Thus, our focus is on improving the factuality of the summaries.
@@ 2.2 Character Knowledge Graph (CKG)
Since characters are integral to narrative Gurung and Lapata (2024), prior work has aimed to construct a graph to easily utilize them. In narrative texts, CKG shows the unidirectional relationship between a subject and an object character. This process is similar to creating a triple (subject-predicate-object) list in knowledge graph construction Chen et al. (2020). Andrus et al. (2022) utilized the OpenIE system Angeli et al. (2015) for story completion and question-answering tasks, integrating it with GPT-3 Brown et al. (2020) to enhance its effectiveness. Alternatively, a recent method Zhao et al. (2024) that assembles CKG directly using LLMs is a more robust approach, as it better captures the nuanced and complex relationships. Our distinction lies not only in constructing CKGs but also in utilizing them to measure and enhance factuality.
@@ 3.1 CKG Extraction
Given a narrative represented as a collection of scenes $\mathcal{N}=\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{m}\}$, where $m$ denotes the number of scenes, the goal is to extract a graph $G$ that encapsulates character relationships. Each scene $\mathcal{S}_{i}\ (1\leq i\leq m)$ is processed individually to extract relation triples (subject-predicate-object) using GPT-4o-mini OpenAI (2023), as detailed in Section E.1. The extracted triples are used to initialize the nodes and determine the edges based on the main relationships between the nodes, forming the final CKG $G$ through the following two steps.
@@
First, to maintain consistency in character identification, we construct a names graph $G_{name}$, consolidating aliases or variations in names in scenes. Our framework processes each scene in turn, extracts all character names, and determines several names refer to the same character based on the context using LLM. For example, in The Lord of the Rings, ‘Frodo’ and ‘Frodo Baggins’ are recognized as the same character. As illustrated by ‘Frodo / Frodo Baggins’ in Figure 3(a), each name variation is a node. [^2] In practice, an undirected edge is added between nodes that refer to the same character. This step ensures an accurate capture of relationships, even when names vary across scenes. The CKG is initialized using names from the names graph.
Second, to preserve the consistency of relationships, we sample extracted triples multiple times Wang et al. (2023); Brown et al. (2024) and select frequent ones as the final edges. Let the node set $V$ be the set of all characters in the names graph:
@@
Only triples with named entities as subjects and objects are used; if an object is missing, a self-loop is added to represent the state of a character. We then define the edge set $E$ of our CKG as
@@
For triples with the same subject and object, frequent predicates capture temporal changes as directed edges. For triples with the same subject and object, our CKG saves predicates that occur frequently across multiple scenes. By storing these frequent predicates in sequential order, our framework captures temporal narrative shifts as directed edges (Eq 2). For example, since the early scenes show that ‘Frodo’ fears ‘Sauron’ and the later scenes show that he resists ‘Sauron’, the CKG in Figure 3(a) displays two distinct relationships. The process of deciding edges is repeated to construct a CKG that can effectively evaluate the factuality of summaries.
@@ 3.2 NarrativeFactScore Calculation
We invent a new metric to guide agentic evaluation, unlike existing factuality metrics Min et al. (2023); Bishop et al. (2024) that do not provide evidence or feedback for their scores, by considering events in the input story superficially but overlooking relational information about characters. Our metric addresses these limitations by incorporating character relationship graphs and providing detailed feedback. To calculate the factuality of the narrative summary, we first generate an initial summary $Z$ using the prompt described in Section E.2.
To evaluate the factuality of the initial summary $Z$, we decompose it into smaller verifiable units, similar to the approach used in Min et al. (2023). Using the prompt in Section E.3, each sentence in the initial summary $Z$ is divided into a list of atomic facts $A=\{a_{1},a_{2},\ldots,a_{z}\}$.
To evaluate each atomic fact $a_{k}$, we need the scene and information about the characters that appear in the atomic fact. First, we retrieve the most relevant scene $\mathcal{S}_{i}$ within the narrative $\mathcal{N}$, by using the BGE-M3 Chen et al. (2024). Second, we also retrieve the subgraph $g$ from the linearized CKG $G$, as illustrated in Figure 3(b), by sorting candidate triples by similarity to the atomic fact $a_{k}$.
@@
where $\mathbbm{1}$ is the indicator function, yielding 1 if the atomic fact $a_{i}$ is factual and 0 otherwise. This evaluation is carried out using the prompt detailed in Section E.4, which produces 1 if the atomic fact is accurate. If the atomic fact is determined to be inaccurate, then feedback $f_{k}$ is provided on how to correct it. Finally, the NarrativeFactScore is calculated as the proportion of atomic facts that are found to be factual, defined by the following equation:
@@ 3.3 Agent-based Fact Refinement
The new metric leveraging consistent CKG enables the LLM agent to guide refinement by using feedback from the evaluation. This process involves three key inputs: original narrative to provide global context, the initial summary that requires modification, and the feedback detailing the inaccuracies and reasons for those errors. Using a prompt that incorporates these inputs in Section E.5, LLM refines the initial summary by correcting only the factually inaccurate parts provided from feedback, then generates the improved summary $y$.
@@ 4.1 Implementation Details
To generate the initial summary $Z$, we adopt hierarchical merging Chang et al. (2024) that ensures the logical structure of the narrative is preserved. The narrative is first divided into chunks $C_{i}$ where each chunk is formed incrementally by adding scenes until a predefined context size444we set the predefined context size of a chunk to 1024. is reached. Once this limit is exceeded, a new chunk begins, resulting in a sequence of chunks $\mathcal{C}=\{C_{1},C_{2},...,C_{n}\}$. Each chunk $C_{i}$ is then independently summarized using the prompt specified in Section E.2, and the resulting chunk summaries are sequentially merged to produce the initial summary $Z$.
@@ 4.3 Correlation with Human Factuality Scores
Dataset # tokens (source) # scenes # tokens (summary)
FABLES 127,467 38 594
STORYSUMM 782 322
Table 1: Dataset statistics for STORYSUMM and FABLES.
@@
To evaluate whether the NarrativeFactScore we proposed correlates effectively with human factuality, we required benchmarks that satisfy two conditions. First, they should provide human factuality scores for multiple LLM-generated summaries of each narrative, enabling correlation analysis. Second, they should be released after October 2023 to minimize potential contamination given the GPT-4o-mini knowledge cutoff. STORYSUMM Subbiah et al. (2024) and FABLES Kim et al. (2024) are the only datasets that meet both conditions, so we chose them as our benchmarks in Tables 6 and 3.
To further explain these datasets, Table 1 shows their statistics. STORYSUMM is relatively short (782 tokens on average, 322 in summaries) and thus scenes were not separated, while FABLES is substantially longer (>100K tokens with 38 scenes). These statistics confirm that our method effectively calculates factuality for both short- and long-length narratives.
@@ 4.4 Summarization Performance Evaluation
We evaluated our framework on the MENSA Saxena and Keller (2024b) and MovieSum Saxena and Keller (2024a) datasets. MENSA aligns movie scenes with Wikipedia summaries and MovieSum pairs screenplays with summaries. Since both datasets provide ground-truth summaries, they can be used to test whether our agent-based refinement improves both factuality and summarization quality. We use the full test sets: 50 samples from MENSA and 200 from MovieSum.
@@ 5.1 How Consistently Does Ours Capture Character Relationships?
To effectively evaluate factuality and improve summary, it is necessary to generate an accurate and consistent CKG. According to Kim et al. (2024); Zhao et al. (2024), the “naive extract” approach, where an LLM extracts character relationships in one step, often fails to consistently capture some relationships. Thus, our objective is to verify whether our approach can generate a consistent CKG. Unlike other datasets, Conan Zhao et al. (2024) provides ground truth annotation of character relationships within narratives. To evaluate whether the generated relation is semantically similar to this ground truth, we measure the BERTScore Devlin et al. (2019).
As shown in Table 5, our method generates CKG that is closely similar to ground truth. Although the “naive extract” achieves 86.26, it occasionally produces incorrect relationships. In contrast, by reasoning about relationships scene by scene and aggregating them, our method chooses more accurate relationships and constructs a consistent CKG. This supports the role of the CKG in enhancing factuality assessment, as shown in Table 3. Even without explicit scene boundaries, as in the Conan dataset, we segmented the text into 256-token chunks with 128-token overlaps. This approach still showed a high BERTScore, demonstrating robustness.
@@ 5.2 Challenging Set
We aim to evaluate whether our metric can provide feedback necessary to improve factuality in recent narratives. Although LLM-based metrics provide accurate factuality feedback for narratives within their pretraining data, they tend to be less reliable for narratives outside of their training corpus. However, our metric provides accurate feedback by evaluating summaries based on narrative story and character relationships rather than relying on parametric knowledge alone. Therefore, we define a challenging set of works published after the knowledge cutoff date of our LLM to verify whether our metric improves factuality through its feedback.